OpenStudy (anonymous):

lim x-->-2^- (x-3)|x+2|/(x+2)

myininaya (myininaya):

|x+2|=-(x+2) since x<-2

myininaya (myininaya):

\[\lim_{x \rightarrow -2^-}-\frac{(x-3)(x+2)}{x+2}=\lim_{x \rightarrow -2^-}-(x-3)=-(-2-3)=-(-5)=5\]

OpenStudy (anonymous):

should i follow the same procedure in case of lim x--> 2+^

myininaya (myininaya):

is that x->-2^+?

OpenStudy (anonymous):

yes

myininaya (myininaya):

that would mean |x+2|=x+2 since x>-2

myininaya (myininaya):

\[\lim_{x \rightarrow -2^+}\frac{(x-3)(x+2)}{x+2}=\lim_{x \rightarrow -2^+}(x-3)=-2-3=-5\]